What is the perimeter and area of a square if the diagonal is 16 meters
Answer by harpazo(655)

A = side^2
Perimeter = 4s
s^2 + s^2 = 16
2s^2 = 16
s^2 = 16/2
s^2 = 8
Take square root of both sides.
s = sqrt(8)

I found the value of one side of the square.

Use the formulas given above for Area and Perimeter to find your answers.